OpenStudy (anonymous):

Given the function y = 5x – 7, how does the y-value change as x increases by 1? A. The y-value decreases by 7. B. The y-value decreases by 2. C. The y-value increases by 12. D. The y-value increases by 5.

OpenStudy (johnweldon1993):

Well lets check...lets use x = 0 y = 5x - 7 y = 5(0) - 7 y = -7 Okay...now increase 'x' by 1...so we'll be using x = 1 y = 5x - 7 y = 5(1) - 7 y = 5 - 7 y = -2 so we went from -7 to -2...what happened?

OpenStudy (anonymous):

decrease by 7

OpenStudy (johnweldon1993):

Not quite....-7 to -2 is a 5 unit increase right?  So y increases by 5

OpenStudy (skullpatrol):